a) A stone is suspended from one end of a thread and execute a simple harmonic motion with a frequency of 0.5 Hz. If the length of the thread is increased by four times the initial length, then determine the period of the harmonic motion. 


1
Expert's answer
2021-05-19T11:03:36-0400

The period of the harmonic motion is given as follows:


"T = 2\pi \sqrt{\dfrac{l}{g}}"

where "l" is the length of the thread, and "g" is the gravitational acceleration.

On the other hand, by definition, the period is the inverse frequency:


"T = \dfrac{1}{f} = \dfrac{1}{0.5Hz} = 2s"

If we increase the length by four times: "l_1= 4l", then the period increases by 2 times:


"T_1 =2\pi \sqrt{\dfrac{l_1}{g}} = 2\pi \sqrt{\dfrac{4l}{g}} = 2\cdot 2\pi \sqrt{\dfrac{l}{g}} = 2T"

Since "T = 2s", the nes period is "T_1 = 2\cdot 2s = 4s".


Answer. 4s.